OpenStudy (anonymous):

Janet has a jar of 54 coins consisting of dimes and quarters. If she has a total of $8.70, how many dimes does she have?

OpenStudy (campbell_st):

you need 2 equations with 2 unknowns let d = dimes and q = quarters then d + q = 54 (1) number of coins 0.1d + 0.25q = 8.70 (2) value of the coins substitution method is probably the easiest rewrite equation (1) q = 54 - d substitute into equation 2 0.1d + 0.25(54 - d) = 8.70 expand and collect like terms 0.1d + 13.5 - 0.25d = 8.70 -0.15d + 13.50 = 8.70 -0.15d = - 4.80 d = 32 the container holds 32 dimes no substitute into equation (1) to find the quarters 32 + q = 54 q = 22 checking 0.1 x 32 + 0.25 x 22 = 3.20 + 5.50 = 8.70